Greetings. I have a word problem that I must use logarithms to figure out.
"On a certain street, a section having an area of 4896.7 square yards is paved at an expense of $12,642,17. What would be the expense to a property holder who pays for 356.28 square yards of the pave?
This is not from a textbook. The questions are similiar to those found in the Penn Foster study units of Technical Math I.

Answer by stanbon(48535)

On a certain street, a section having an area of 4896.7 square yards is paved at an expense of $12,642,17. What would be the expense to a property holder who pays for 356.28 square yards of the pave?
-----------------------------
Let "x" be the cost for the 356.28 sq. yds.
12,642.17/4896.7 = x/356.28
2.58177344 = x/356.28
x=$919.83
Cheers,
Stan H.